Find the slope of each line
y=-x/4 + 8
i plug in the 0 and have 0/4 +8
whats next

Answer by lwsshak3(11628)

Find the slope of each line
y=-x/4 + 8
**
This is an equation of a straight line.
Its standard form: y=mx+b, m=slope, b=y-intercept
For given equation:y=-x/4 + 8
slope=-1/4